OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! –5(1 – 5x) + 5(–8x – 2) = –4x – 8x A.x = –5 B.x = 0.19 C.x = 0.55 D.x = 5

OpenStudy (anonymous):

A, x=-5

OpenStudy (jmark):

–5(1 – 5x) + 5(–8x – 2) = –4x – 8x -5+25x-40x-10 = -–4x – 8x -15x-15 = -12x -3x = 15 x = -5 A . x = -5 ICSE maths question paper for class 9 - http://icse.edurite.com/icse-question-papers/icse-maths-question-papers-for-class-9.html